OpenStudy (anonymous):

If dy/dx=20/(x+1)^3 Find d^2y/dx^2 and the real values of x for which d^2y/dx^2 = -15/4

OpenStudy (anonymous):

\[\frac{dy}{dx}=\frac{20}{(x+1)^3}=20(x+1)^{-3}\\ \frac{d^2y}{dx^2}=-60(x+1)^{-4}=-\frac{60}{(x+1)^4}\] Now you solve for x in the following equation: \[-\frac{60}{(x+1)^4}=-\frac{15}{4}\]

OpenStudy (anonymous):

So you rearrange \[-\frac{ 60 }{(x+1)^4 } =-\frac{ 15 }{ 4 }\] with respect to x. does that give \[x=\sqrt[4]{\frac{ 15 }{ 4\times-60 }}-1\] ? if so I get a math error

OpenStudy (anonymous):

\[\begin{align*} -\frac{60}{(x+1)^4}&=-\frac{15}{4}\\ \frac{60}{(x+1)^4}&=\frac{15}{4}\\ \frac{4}{(x+1)^4}&=\frac{1}{4}\\ \frac{16}{(x+1)^4}&=1\\ 16&=(x+1)^4\\ \pm\sqrt[4]{16}&=x+1\\ \pm2&=x+1\\ x=1~&~x=-3 \end{align*}\]

OpenStudy (anonymous):

x is equal to 1 or -3

OpenStudy (anonymous):

Ah ok my mistake I read value of x instead of values, thanks for the help